OpenStudy (anonymous):

Simplify by removing factors of 1. (5r^2+5r)/(20r^2+25r)

OpenStudy (unklerhaukus):

\[\frac{5r^2+5r}{20r^2+25r}\]\[=\frac{5r(r+1)}{5r(4r+5)}\]\[=\cdots\]

OpenStudy (anonymous):

???

OpenStudy (unklerhaukus):

now just cancel common factor in the numerator and denominator

OpenStudy (anonymous):

and how do you do that?

OpenStudy (unklerhaukus):

\[=\frac{\cancel{5r}(r+1)}{\cancel{5r}(4r+5)}\]\[=\cdots\]

OpenStudy (anonymous):

The simplified form is?